Sol Summary Report – March 5th

Crew 275 Sol Summary Report 05Mar2023

Sol: 21

Summary Title: Sunday Slumber

Author’s name: Jeremy Rabineau, Commander.

Mission Status: Nominal.

Sol Activity Summary:

With the high winds, we decided to limit our EVA to changing the batteries and retrieving the data of the atmospheric instruments. The weather bothered the two EVA crewmembers, but they still managed to conduct all the operations required. Onsite, they noticed that the LOAC and the Mega-Ares were both off. Apparently, the LOAC stopped working after being turned on during the last EVA, while the Mega-Ares worked for about 40 hours before it was out of battery. We are currently inspecting what may have caused the dysfunction of the LOAC.

It was definitely a strange Sunday. While they are usually dedicated to getting some rest and cleaning the Hab, this time we had to get back on track with all the activities that have been delayed due to the medical event of Friday night. The good point is that the station had already been fully cleaned prior to the arrival of the journalists on Friday.

Unfortunately, the weather was also too windy for astronomy activities to be conducted. The afternoon was then dedicated to other activities: some crewmembers watched a movie, while others did some physical activity, and others took a nap. We finally took some time to do some data processing and to train with AI4U, an AI tool aimed at assisting the astronauts in their daily tasks.

Look Ahead Plan:

All the EVAs involving the injured crewmember have been cancelled until the end of the simulation. The optimistic plan is to see the surgeon as early as tomorrow (Sol 22), so that they can potentially have surgery on Tuesday (Sol 23). We will know more about this plan when we can call the surgeon tomorrow at 9:00. We will make sure to keep close contact with Mission Support, to communicate all the information that may be relevant to the health of this crewmember as well as to the continuation of the simulation. So far, we plan to use the CrewCar to proceed with the transport of the injured crewmember.

Tomorrow, we are also planning to do the photogrammetry EVA that had to be postponed following the medical event of Friday evening. We will take this opportunity to change the batteries of the atmospheric instruments, since they apparently have problems lasting even 48 hours.

A media visit is planned for next Tuesday (Sol 23) with a crew from France Televisions. They will arrive at 9am and depart at 5pm.

Anomalies in work: None.

Weather: Windy, low 2°C/36°F, high 12°C/54°F.

Crew Physical Status: One crewmember injured (broken collarbone / clavicula), awaiting appointment with a surgeon early next week (Sol 22) in Richfield.

EVA: EVA # 17 was successfully performed, even though some anomalies were observed onsite.

Reports to be filed: Sol Summary, EVA Report, EVA Request, Operations Report, Journalist Report, GreenHab Report, HSO Report.

Support Requested: None.

Copyright © The Mars Society. All rights reserved. | Main Site